OpenStudy (1645323):

The dimensions of a rectangular prism are length = (x -1), width = (x - 5), and height = (x - 7). Find the approximate relative maximum value for the volume of the rectangular prism. Select one: a. 14.2 b. 16.9 c. 18.5 d. 10.3

OpenStudy (1645323):

@ganeshie8 @phi

OpenStudy (phi):

You should complain to your teacher about this question. They want you to either use calculus or plot the curve and look for a local maximum If we plot the curve we get the attached graph. *However*, notice the local maximum occurs at approximately x= 2.57 this means the dimensions of your prism will have negative values. Example, width = (x - 5) = -2.43 which is clearly bogus. They should just ask, what is the local maximum for the cubic curve defined by y= (x-1)(x-5)(x-7) when they introduce the idea that these are the dimensions of a prism they went off the tracks (and should not be allowed to make up math questions)